# Self-supervised Learning for Geospatial AI: A Survey

### v1

@@ IV-B3 Applications
Similar to the representations derived for road networks, trajectory representations can be directly utilized or fine-tuned for downstream applications. These representations, regardless of their specific context information, facilitate a variety of operations due to their vectorized format, including similarity computation [97, 133] and clustering [103, 95]. Moreover, they can be utilized to support diverse applications, such as transportation mode prediction [100], driver status inference [104], and anomalous trajectory detection [98]. In scenarios involving map-matched trajectories, these representations prove particularly valuable in applications focused on path analysis, such as travel time estimation [119], path ranking [120], and destination prediction [117]. In addition, semantic trajectory representations trained with SSL methods are typically fine-tuned to enhance performance in next location prediction [121], and can be uniquely applied in trajectory user-linking task [124, 123]. These applications demonstrate the broad utility and adaptability of SSL for trajectories in advancing the understanding and analysis of mobility patterns and intelligent transportation systems.
V Polygons
A wide range of geospatial objects are often represented with polygonal geometries. This type provides a detailed depiction of the shapes of geospatial objects in 2D planes. In principle, from a 2D perspective, many other geospatial objects discussed in this paper, such as POIs and road segments, can also be represented as polygons. However, due to the substantial effort required to collect and curate objects with precise polygonal geometries, only a few geospatial data types are represented this way. Within the scope of this survey, a predominant object type that has garnered significant interest in an SSL context is the modeling of small urban areas, also known as urban regions.
Urban regions have been an ideal analytical scale for surveying and predicting various socio-economic and demographic characteristics in our cities [134]. These regions are essentially small polygonal areas partitioned by specific methods, such as fine-grained areas defined by road networks, fixed-size grids, or administrative units like Singapore Subzones [135] and NYC census Tracts [136]. Regardless of the partitioning method, modeling urban regions primarily relies on other source data modalities that reflect the intrinsic or contextual properties of these regions. Thus, a region can be viewed as a “container” of various human activities and data entities from different modalities. Some data modalities, such as POIs and street view images (SVIs), capture the intrinsic properties of the regions, while others, like human trajectories, describe region traits from a connectivity perspective, illustrating how people traverse through regions. In addition, spatial proximity is an important contextual factor, as spatially closer regions tend to be more similar.
With such diverse information perspectives used to model urban regions in an SSL context, most studies incorporate multiple perspectives. For example, a strand of region embedding studies [137, 138, 139, 140, 141, 142, 143, 144] utilizes both POIs and human trajectories, sometimes supplemented by the consideration of spatial proximity. The modeling methods of different perspectives (data modalities) are often intertwined and challenging to disentangle. In this section, our goal is to present each study holistically, rather than fragmenting the analysis into fragmented components. This means that if a study models certain contextual information, the entire analytical pipeline of that study will be discussed collectively. The surveyed studies for urban regions are listed in Table IV.
@@ V Polygons
TABLE IV: A summary of the surveyed papers on self-supervised learning for urban regions.
@@
Reconstruction of node feature
@@
Human mobility InfoNCE Region-region contrast
@@
Region-region contrast
HAFusion[144] Predictive
Human trajectory, POI,
Land use
None Human mobility MSE, KL divergence
Reconstruction of human
mobility patterns, POI,
and land use feature similarity
@@
Reconstruction of images,
and human mobility patterns ,
Multi-view graph discrimination
@@
Text autoregressive prediction
Image- and token-level
Image-text alignment
@@
Point injection, Building removal
Spatial proximity InfoNCE, Triplet loss Region-region contrast
V-A Urban regions
V-A1 Intrinsic Attributes
Among various data modalities, POIs and imagery data, including remote sensing (RS) images and SVIs, are most commonly utilized to reflect the intrinsic properties of each region. POIs capture socioeconomic factors, while images reflect the overall physical appearance from a human or aerial perspective. Several studies have explored these data sources to learn region representations in a self-supervised manner, sometimes with minimal consideration of contextual information.
V-A2 Context information
POIs/knowledge graphs. Based on the pioneering work [145], SKRL4RS [146] further enhances the information of spatial entities by incorporating the rich context information from two well-established knowledge graphs, YAGO and DBpedia. Instead of one-hot category encoding, the spatial entities in the knowledge graphs are transformed into representations that encapsulate the relatedness derived from their hierarchical categories (ontologies) and their proximity within the knowledge graph. In this way, the semantics of spatial entities within regions are better captured. For example, the similarity between a Japanese restaurant and a Korean restaurant is much larger than that between a factory and a Korean restaurant. In another research line, HGI [147] extends beyond the POI category embedding technique in [64] by further applying GNN aggregation process to escalate the representations to region and city levels. This model is optimized by maximizing the mutual information among the POI-region-city hierarchy.
@@ V-A2 Context information
POIs + Human trajectories (+others). A large portion of region representation learning studies incorporate POIs alongside human trajectories, as these two data sources naturally complement each other. POIs outline the range of human activities within each region, while human trajectories illustrate the interconnections and relatedness between regions. This integration is sometimes augmented with additional data sources, such as land use data, to further enhance the effectiveness of the learned region representations. In this regard, MP-VN [137] constructs two POI graphs to capture both static and mobility patterns among various POI types, based on their mobility connectivity patterns and average geographic distances. The two graphs are flattened as input to an autoencoder designed to learn region representations through graph reconstruction. Besides, the model incorporates the inter-region relations considering both spatial proximity and functional similarity derived from POIs to enhance the learning process. Built on this, DLCL [138] employs an adversarial autoencoder to learn representations of the two POI graphs in [137]. CGAL [139] further enhances the model by introducing a collective adversarial learning approach instead of autoencoder reconstruction. Specifically, it implements an assemble-disassemble strategy in which the assemble step fuses the two graph views into region representations, and the disassemble step then disaggregates the fused representations to reconstruct the original two graph views. It also incorporates two adversarial components to capture the similarities between regions based on POI distributions, textual information, and temporal patterns reflected from human trajectories. This model is trained to minimize the reconstruction loss of the adversarial framework. HAFusion [144] utilizes human trajectories, POIs, and land use data to model urban regions. Each region is characterized by a human mobility feature derived from outflow records, a POI category feature, and a land use feature quantifying the counts of land use zones in different categories. This method employs an attention-based encoder to capture region correlations within a single view and across different views. Furthermore, it applies another attention-based fusion module that integrates multi-view embeddings and captures higher-order correlations among regions.
@@
Another line of research extensively emoplys GNNs. MVURE [152] utilizes human trajectories, POIs, and user check-ins to construct four different graph views. It then employs graph attention networks to encode each view and implements cross-view information sharing with attention mechanisms as well as multi-view fusion using adaptively weighted combination of different views. The model is trained through objectives of reconstructing human mobility patterns and region relatedness reflected by POIs and check-ins. HREP [140] further considers several types of graph edges using human mobility data, POIs, and geographic context. This method constructs source and target edges based on human trajectories, POI edges derived from the similarities of POIs shared by different regions, and geographic context edges based on spatial adjacency. Using this heterogeneous graph structure, region representations are learned through attention-based GNN method by three objectives: a geographic proximity loss, a mobility reconstruction loss, and a POI correlation reconstruction loss. Additionally, HREP employs prompt learning (prefix-tuning) to adapt the learned representations for various downstream urban analytical tasks, a technique adapted from natural language processing. In addition, HUGAT [153] utilizes POI check-in trajectories and land use data for region representation learning. It defines five types of nodes, such regions and POI categories, and two types of edges representing spatial and temporal relations. This method then constructs a heterogeneous information network and designs five meta-paths to involve interesting relationships between regions, such as identifying regions that are popular destinations at the same time. A heterogeneous graph attention network [163] is employed to derive region representations by minimizing the difference between estimated and actual mobility patterns of regions, check-in distributions, and land use distribution. Region2Vec [141] utilizes inter-region human trajectories, spatial adjacency, and POI distributions as three similarity measures to construct a multi-graph for regions. Specifically, human mobility similarity is reflected by the co-occurrence frequency between regions, and POI distribution similarity is measured using the embeddings from a POI knowledge graph. This method utilizes GNN and fusion layers based on attention mechanisms to generate comprehensive region representations, with the training objectives of reconstructing human mobility patterns, preserving geospatial adjacency, and POI distribution similarity.
@@
The third line of research employing POIs and human trajectories focuses on the application of contrastive learning paradigm. ReMVC [142] is a dual-view approach integrating POIs and human mobility data. For the POI view, it uses region-level POI category proportions as raw features, and performs data augmentation through random POI insertion, deletion, and replacement. These augmented POI views of regions serve as positive samples, whereas POI data from differing regions are used as negative samples for the contrastive learning. For the human mobility view, it constructs two heatmaps to represent the source and destination patterns of each region, followed by the augmentation of Gaussian noise injection to form positive samples for the contrastive learning. Furthermore, an inter-view contrastive learning objective is employed to ensure that representations of a region from different perspectives are similar, while the representations of different regions are distinguishable. ReCP [143] develops a fusion technique for multiple information views from an information theory perspective. Apart from maximizing the mutual information (consistency) shared between different views, this method implements a dual prediction strategy to minimize the conditional entropy between representations from different views, thereby reducing the inconsistency between views.
@@
Tile2Vec [154] generates representations based on RS images patches for square-shaped regions. It employs a CNN model to encode remote sensing images and a triplet loss to ensure that patches which are geographically adjacent are also close in the embedding space. RegionEncoder [155] leverages RS images, POIs, and human trajectories for learning region representations. Initially, it employs a denoising autoencoder to extract representations from RS images. Following this, a region graph is constructed, utilizing region-level POI features as node attributes and human trajectory data to define inter-region connectivity through edges. The model is trained with three specific objectives: a reconstruction loss for the denoising autoencoder, a reconstruction loss for human mobility patterns, and a binary cross-entropy loss used by a discriminator to verify whether RS image representations and region graph node representations correspond to the same geographical region. The study in  [158] adopts a contrastive learning strategy to maximize the similarity of representations derived from RS images that are geographically adjacent and those that have similar POI distributions. The representations learned from the two pathways, including spatial proximity and POI distribution similarity, of each region are adaptively fused using learnable weights in downstream tasks. Urban2Vec [156] derives the initial region representations by averaging all the SVI representations obtained from a CNN-based model for each region. It subsequently fine-tunes this image encoder with a spatial proximity-based triplet loss to bring SVIs that are spatially close together in the embedding space. Besides, this method integrates the POIs by further generating the POI representation of a region that encapsulate all the words associated with the POIs in each region. It introduces another triplet loss designed to merge POI information into the region representations by minimizing the distance between each region’s representation and its corresponding POI representation. M3G [157] extends Urban2Vec by simultaneously utilizing SVIs, POIs, and human trajectories. The techniques for handling SVIs and POIs are similar to those in Urban2Vec. Additionally, M3G enhances the model by conducting region-level contrastive learning, selecting other regions that are either spatially close or connected through human mobility as positive samples. The model is trained using a combination of triplet losses: region-SVI, region-POI, and region-region, which consider both spatial proximity and human mobility connections.
@@
OpenStreetMap data. A new trend in region representation is to utilize data from OpenStreetMap (OSM) [67], a global-scale open geospatial dataset contributed by a community of mappers. OSM offers a extensive repository of geospatial entities such as building footprints and road networks, serving as valuable and accessible resources for learning effective region representations. RegionDCL [162] extracts building footprints and POIs from OSM, and begins by encoding the shape information from building footprints using a CNN-based model to generate initial region features. Furthermore, RegionDCL addresses the challenge of empty areas, which are spaces not explicitly represented in discrete geospatial vector data but are physically present. To effectively represent these empty areas in the final region embeddings, this method employs Poisson Disk Sampling to fill these gaps. Each building group–small regions partitioned by road networks–is then processed through a distance-biased Transformer. This Transformer is trained using building group-level contrastive learning. Subsequently, these building groups are aggregated into larger regions for a second round of contrastive learning, employing a triplet loss with an adaptive margin to refine the final region embeddings.
V-A3 Applications
@@
As urban regions serve as a critical analytical scale for various urban analyses and prediction tasks, the learned region representations are used in a diverse range of downstream tasks. Commonly, these region embeddings are utilized by integrating them as frozen inputs into shallow task predictors, such as MLP, for making task-specific inferences. Additionally, several studies have enhanced the utilization of region representations in downstream tasks through advanced methods like prompt learning [140] and adaptive multi-view fusion [158]. From the perspective of downstream tasks, the predominant tasks involve predicting various socioeconomic indicators in cities. These include region-level attributes such as land use/urban function/land cover [149, 142, 143], population density [156, 147], house prices [155, 160], average income [157, 148], check-in counts [137, 139], crime rates [151, 148], service call volumes [144], GDP [159, 160], nighttime lighting [161], takeaway order volumes [158], health indices [154], and so on. Moreover, region representations are extensively utilized for tasks like similar region search[145, 146] and region or land use clustering [141] in a fully unsupervised manner, which has significant practical implications for real-world urban planning and management. Furthermore, region representations can benefit studies on human mobility in cities, such as predicting human flow and bike flow [150, 153].
The application of SSL-based region representations has proven to be diverse and effective across many critical urban analytical tasks, leading to substantial real-world benefits. Such applications not only improve the accuracy of urban analyses but also support more informed decision-making in urban planning and policy-making. For example, accurate predictions of socioeconomic indicators can inform better resource allocation, enhance public services, and facilitate targeted interventions in sectors like healthcare, education, and transportation. Additionally, the capability to identify and cluster similar regions supports strategic site selection, the development of cohesive urban zones, optimization of land use, and promotion of sustainable development. As urban environments continue to expand and evolve, the role of robust region representations in tackling complex urban challenges will become increasingly crucial, paving the way for smarter, more resilient cities.
VI Discussion
Sections III-V have presented specialized SSL techniques designed for each geospatial data type. In this section, we extend the discussion by presenting several aspects not covered in the preceding sections. Specifically, we present some emerging trends in SSL techniques for geosaptial AI, including multi-type learning, and the adaptation to textual space via LLMs. Moreover, we provide a brief introduction of research that adopts task-specific SSL for geospatial applications.
VI-A Multi-type Learning

### v2

@@ 4.2.3 Applications
Trajectory representations can be directly utilized or fine-tuned for downstream applications. These representations, regardless of their specific context information, facilitate a variety of operations due to their vectorized format, including similarity computation [92, 180] and clustering [96, 91]. Moreover, they can be utilized to support diverse applications, such as transportation mode prediction [94], driver status inference [97], and anomalous trajectory detection [93]. In scenarios involving map-matched trajectories, these representations prove particularly valuable in applications focused on path analysis, such as travel time estimation [114], path ranking [115], and destination prediction [112]. Semantic trajectory representations trained with SSL methods are typically fine-tuned to enhance performance in next location prediction [117], and can be uniquely applied in trajectory user-linking task [121, 119]. These applications demonstrate the broad utility and adaptability of SSL for trajectories in advancing the understanding and analysis of mobility patterns and intelligent transportation systems.
5 Polygons
Polygonal representations offer detailed 2D shapes of geospatial objects. In principle, many geospatial objects like POIs and road segments could be represented this way, while the effort to collect precise polygonal data limits its use. In this survey, we focus on urban regions, i.e., small urban areas which serve as the analytical units for a wide array of urban analytical tasks, e.g., for land use, population density, and house price.
Urban regions are partitioned, e.g., by road networks, grids, or administrative boundaries like Singapore Subzones [181] and NYC Census Tracts [182]. These regions serve as ”containers” for data in various modalities, entailing either intrinsic properties of regions, or interrelatedness with other regions. Different types of spatial piror, e.g., spatial proximity, is also crucial, to incorporate the inherent relations entailed by geographic locations. Most studies utilize multiple data perspectives, e.g., the region embedding studies [136, 137, 138, 144, 141, 143, 145, 139] use POIs, human trajectories, and spatial proximity. The modeling of different perspectives is often intertwined, so we present each study holistically, covering the entire analytical pipeline. To this end, we categorize the studies according to the data modalities used. See Table LABEL:tab:_region for surveyed studies.
5.1 Urban regions
5.1.1 Intrinsic Attributes
Among various data modalities, POIs and imagery data, including remote sensing (RS) images and street view images (SVIs), are most commonly utilized to reflect the intrinsic properties of each region. POIs capture socioeconomic factors, while images reflect the overall physical appearance from a human or aerial perspective. Several studies have explored these data sources to learn region representations in a self-supervised manner, sometimes without much consideration of contextual information.
5.1.2 Context information
POIs/Knowledge Graphs. Based on the pioneering work [126], SKRL4RS [131] further enhances the information of spatial entities by incorporating the rich context information from two well-established knowledge graphs, YAGO and DBpedia. Instead of one-hot category encoding, the spatial entities in the knowledge graphs are transformed into representations that encapsulate the relatedness derived from their hierarchical categories (ontologies) and their proximity within the knowledge graph. In this way, the semantics of spatial entities within regions are better captured. For example, the similarity between a Japanese restaurant and a Korean restaurant is much larger than that between a factory and a Korean restaurant. In another research line, HGI [130] extends beyond the POI category
@@ 5.1.2 Context information
Table 4: A summary of the surveyed papers on self-supervised learning for urban regions.
@@
Reconstruction of
node feature
@@
HAFusion[139] Predictive
Human
trajectory,
POI,
Land use
None
Human mobility,
POIs
MSE, KL divergence
Reconstruction of human
mobility patterns, POI,
and land use feature
similarity
@@
Human mobility,
POIs
InfoNCE Region-region contrast
@@
Region-region contrast,
Region feature reconstruction
@@
Reconstruction of images
and human mobility
patterns , Multi-view
graph discrimination
@@
Image- and token-level
Image-text alignment
@@
Text autoregressive
prediction
GeoHG[151] Contrastive RS, POI Heterogenous graph Spatial adjacency, Region feature similarity InfoNCE Graph contrast
ReFound[153]
Predictive,
Contrastive
RS, POI
Foundation model distillation
and text generation
None
KL divergence
Cross-entropy
Knowledge distillation
Masked data modeling
Cross-modal spatial alignment
MuseCL[152] Contrastive RS, SVI, POI None Region feature similarity
Triplet
InfoNCE
Image-level contrast
Image-text contrast
GeoHG[151] Contrastive RS, POI Heterogenous graph Spatial adjacency, region feature similarity InfoNCE Graph contrast
@@
Point injection,
Building removal
Spatial proximity
InfoNCE,
Triplet loss
Region-region contrast
embedding technique in [54] by further applying GNN aggregation process to escalate the representations to region and city levels. This model is optimized by maximizing the mutual information among the POI-region-city hierarchy.
@@
POIs + Human Trajectories (+others). Many region representation studies integrate POIs with human trajectories, as these data sources naturally complement each other. POIs describe the range of activities within a region, while human trajectories reveal inter-regional connections. This integration is often enhanced with additional data, such as land use. For instance, MP-VN [136] constructs two POI graphs to capture both static and mobility patterns among POI types, based on mobility connectivity and geographic distances. These graphs are sent to an autoencoder for learning region representations via graph reconstruction, while also incorporating spatial proximity and functional similarity derived from POIs. Building on this, DLCL [137] employs an adversarial autoencoder to learn from these two POI graphs. CGAL [138] enhances this model by introducing collective adversarial learning, using an assemble-disassemble strategy where fused region representations are disaggregated to reconstruct the original graphs. It captures region similarities based on POI distributions, textual information, and temporal patterns from human trajectories. HAFusion [139] further utilizes human trajectories, POIs, and land use data to model urban regions. Each region is characterized by mobility features, POI categories, and land use counts, with an attention-based encoder capturing intra- and inter-view region correlations, while a fusion module integrates multi-view embeddings and captures higher-order correlations.
@@
Another line of research extensively employs GNNs. MVURE  [140] uses human trajectories, POIs, and check-ins to construct four graph views, encoding each view with a graph attention network. It implements cross-view information sharing through attention mechanisms and fuses the views using adaptive weighting, aiming to reconstruct mobility patterns and region relatedness from POIs and check-ins. HREP [144] enhances this by defining multiple edge types from human mobility, POIs, and geographic context, with source and target edges based on trajectories, POI similarities, and geographic adjacency. It learns region representations using an attention-based GNN and three objectives: geographic proximity loss, mobility reconstruction, and POI correlation reconstruction. Additionally, HREP uses prompt learning (prefix-tuning) to adapt learned representations for downstream tasks, a method borrowed from NLP. In addition, HUGAT [142] utilizes POI check-in trajectories and land use data for region representation learning. It defines five types of nodes, e.g., regions and POI categories, and two types of edges for spatial and temporal relations. This method then constructs a heterogeneous information network and designs five meta-paths for interesting relations like regions that are popular destinations at the same time. A heterogeneous graph attention network is employed to derive region representations by minimizing the difference between estimated and actual mobility patterns of regions, check-in distributions, and land use distribution. Region2Vec [141] constructs a multi-graph using human trajectories, spatial adjacency, and POI distribution embeddings from a knowledge graph. GNN and attention-based fusion layers are employed to create region representations, with objectives of reconstructing mobility patterns, maintaining spatial adjacency, and preserving POI distribution similarities.
@@
The third line employing POIs and human trajectories focuses on the application of the contrastive learning paradigm. ReMVC [143] is a dual-view approach integrating POIs and human mobility data. For the POI view, it uses region-level POI category proportions as raw features, and performs data augmentation through random POI insertion, deletion, and replacement. These augmented POI views of regions serve as positive samples, whereas POI data from differing regions are used as negative samples for the contrastive learning. For the human mobility view, it constructs two heatmaps to represent the source and destination patterns of each region, followed by the augmentation of Gaussian noise injection to form positive samples for the contrastive learning. Furthermore, an inter-view contrastive learning objective is employed to ensure that representations of a region from different perspectives are similar, while the representations of different regions are distinguishable. ReCP [145] develops a fusion technique for multiple information views from an information theory perspective. Apart from maximizing the mutual information (consistency) shared between different views, this method implements a dual prediction strategy to minimize the conditional entropy between representations from different views, thereby reducing the inconsistency between views.
@@
Tile2Vec [147] generates representations based on RS images patches for square-shaped regions. It employs a CNN model to encode remote sensing images and a triplet loss to ensure that patches which are geographically adjacent are also close in the embedding space. RegionEncoder [146] leverages RS images, POIs, and human trajectories for learning region representations. Initially, it employs a denoising autoencoder to extract representations from RS images. Following this, a region graph is constructed, utilizing region-level POI features as node attributes and human trajectory data to define inter-region connectivity through edges. The model is trained with two reconstruction losses, and a cross-modal alignment objective. The study in  [150] adopts a contrastive learning strategy to maximize the similarity of representations derived from adjacent RS images and those that have similar POI distributions. The representations learned from the two pathways of each region are adaptively fused using learnable weights in downstream tasks.
Urban2Vec [148] derives the initial region representations by averaging all the SVI representations obtained from an image encoder for each region. It subsequently fine-tunes this image encoder with a spatial proximity-based triplet loss to bring SVIs that are spatially close together in the embedding space. Besides, this method integrates the POIs by further generating the POI representation of a region that encapsulate all the words associated with the POIs in each region. It introduces another triplet loss designed to merge POI information into the region representations by minimizing the distance between each region’s representation and its corresponding POI representation. M3G [149] extends Urban2Vec by simultaneously utilizing SVIs, POIs, and human trajectories. M3G enhances the model by conducting region-level contrastive learning, selecting other regions that are either spatially close or connected through human mobility as positive samples. In another work, GeoHG [151] used the semantic segmentation features of the European Space Agency images system for processing remote sensing images, and constructed a heterogeneous graph to capture the spatial adjacency and similarities at the region level reflected from the similarities in environmental and socioeconomic characteristics. The graph model was pretrained using contrastive learning.
Another interesting region embedding study using both street view and satellite images, in conjunction with POIs and human trajectories is MuseCL [152]. The method carries out contrastive learning at multiple levels. First, contrast was conducted for street view images based on travel pattern similarity reflected form human trajectories. Likewise, contrast was carried out for remote sensing images based on POI similarity. The rationale of such pairing is that street view and mobility data entail human movement patterns, while remote sensing and POI data both capture the built environment and land use. The final region embeddings are obtained through fusing street view and remote sensing features, and supplemented with textual information from POIs.
With the rapid advancements in vision foundation models that are pretrained with vast amounts of images or image-text pairs, leveraging the power of foundation models for region embedding has become a trend. To this end,  [153] proposed ReFound with POIs and remote sensing images. One major argument here is the distillation of knowledge contained in foundation models is beneficial. To this end, the POI embeddings are enforced to be close to the embeddings of LLM-generated text, image embeddings are pushed towards the representations generated by vision foundation models, and the similarities between image-POI pairs also imitate the similarities derived from a vision-language model. ReFound is finally pretrained by combining distillation objectives, masked data modeling, and cross-modal modality alignment.
@@
OpenStreetMap. A new trend in region representation is to utilize data from OpenStreetMap (OSM) [159], a global-scale open geospatial dataset contributed by a community of mappers. OSM offers a extensive repository of geospatial entities such as building footprints and road networks, serving as valuable and accessible resources for learning effective region representations. RegionDCL [154] extracts building footprints and POIs from OSM, and begins by encoding the shape information from building footprints using a CNN-based model to generate initial region features. Furthermore, RegionDCL addresses the challenge of empty areas, which are spaces not explicitly represented in discrete geospatial vector data but are physically present. To effectively represent these empty areas in the final region embeddings, this method employs Poisson Disk Sampling to fill these gaps. Each building group–small regions partitioned by road networks–is then processed through a distance-biased Transformer. This Transformer is trained using building group-level contrastive learning. Subsequently, these building groups are aggregated into larger regions for a second round of contrastive learning, employing a triplet loss with an adaptive margin to refine the final region embeddings.
5.1.3 Applications
@@
As urban regions serve as a critical analytical scale for various urban analyses and prediction tasks, the learned region representations are used in a diverse range of downstream tasks. Commonly, these region embeddings are utilized by integrating them as frozen inputs into shallow task predictors, such as MLP, for making task-specific inferences. Additionally, several studies have enhanced the utilization of region representations in downstream tasks through advanced methods like prompt learning [144] and adaptive multi-view fusion [150]. From the perspective of downstream tasks, the predominant tasks involve predicting various socioeconomic indicators in cities. These include region-level attributes such as land use/urban function/land cover [133, 143, 145], population density [148, 130], house prices [146, 129], average income [149, 132], check-in counts [136, 138], crime rates [135, 132], service call volumes [139], GDP [128, 129], nighttime lighting [127], takeaway order volumes [150], health indices [147], and so on. Moreover, region representations are extensively utilized for tasks like similar region search[126, 131] and region or land use clustering [141] in a fully unsupervised manner, which has significant practical implications for real-world urban planning and management. Furthermore, region representations can benefit studies on human mobility in cities, such as predicting human flow and bike flow [134, 142]. Overall, the application of SSL-based region representations has proven to be diverse and effective across many critical urban analytical tasks, leading to substantial real-world benefits. As urban environments continue to expand and evolve, the role of robust region representations in tackling complex urban challenges will become increasingly crucial, paving the way for smarter, more resilient cities.
6 Multi-type Learning